# arXiv:2602.23520v1[cs.IT] 26 Feb 2026

### v1

@@ Abstract
We introduce the zero-incoherence capacity for interactive multi-location encoding systems: the maximum encoding rate that guarantees exactly zero probability of disagreement among replicated encodings. Our main information-theoretic results are compact and self-contained: an exact capacity theorem ($C_{0}=1$), a tight side-information lower bound for resolution ($\geq\log_{2}k$ bits for $k$-way incoherence), and a rate–complexity separation (modification cost $O(1)$ at capacity vs $\Omega(n)$ above).
The paper frames encoding locations as terminals in a multi-terminal source-coding model. Derivation (automatic deterministic dependence) is interpreted as perfect correlation that reduces effective rate; only complete derivation (one independent source) achieves zero incoherence. We give concise achievability and converse proofs in IT style, formalize the confusability/incoherence graph connection, and present an explicit mutual-information argument for the side-information bound.
Theoretical contributions are supplemented by constructive instantiations (programming-language patterns and a software case study). For TIT submission we move detailed language evaluation, extended code examples, and the full Lean proof corpus to supplementary material; the main text contains brief instantiations only. Core theorems (capacity, realizability, bounds) are machine-checked in Lean 4; entropy arguments apply standard Fano-inequality techniques.
Index Terms–zero-error capacity, multi-terminal source coding, side information, mutual information, confusability graph
@@ I Introduction
I-A Zero-Incoherence Capacity
We address this information-theoretic question: what encoding rate guarantees exactly zero probability of disagreement among replicated encodings in an interactive setting? An encoding system stores a fact $F$ (value in $\mathcal{V}_{F}$) at locations $\{L_{1},\dots,L_{n}\}$. The system is coherent when all locations agree; otherwise it is incoherent. The central quantity is the zero-incoherence capacity $C_{0}$, the supremal encoding rate that forces incoherence probability to be exactly zero. Our main theorem is compact:
@@ I-B The Capacity Theorem
I-C Applications Across Domains
@@
I-D Connection to Classical Information Theory
@@ I-E Encoder Realizability
I-F Paper Organization
Core theorems (capacity, realizability, complexity bounds) are machine-checked in Lean 4 [15] (3378 lines, 176 theorem/lemma statements, 0 sorry placeholders). The entropy and mutual-information arguments in Section VI-D are encoded as explicit classical-information assumptions in the Lean closure layer, matching paper-level conditional usage of Fano-style bounds [13].
@@ I-G Core Theorems
Uniqueness. $C_{0}=1$ is the unique capacity: DOF =0=0 fails to encode; DOF >1>1 exceeds capacity. Given zero-incoherence as a constraint, the rate is mathematically forced.
I-H Scope
@@ 1 Coherence guaranteed (optimal rate under coherence constraint)
This is formalized as the Zero-Incoherence Capacity theorem (Theorem II.36) in Section II-K.
@@ III-D Derivation: The Coherence Mechanism
Structural facts require definition-time derivation. Structural facts (class existence, schema structure, service topology) are fixed when defined. Compile-time derivation that runs before the definition is fixed is too early (the declarative source is not yet fixed). Runtime is too late (structure already immutable). Definition-time is the unique opportunity for structural derivation (REQ1, REQ2).
@@ IV-H The Logical Chain (Summary)
Mechanization status. The core requirement chain is machine-checked in Lean 4; current build statistics are 3378 lines, 176 theorem/lemma statements, and 0 sorry placeholders.

### v2

# arXiv:2602.23520v2[cs.IT] 02 Mar 2026

@@ Abstract
Single-Source Coherence Theorem. We prove that among all possible degrees of freedom (DOF = number of independent encoding locations), exactly one value (DOF = 1) guarantees coherence. DOF = 0 fails (no fact encoded). DOF $\geq$ 2 fails (permits explicit construction of inconsistency). Only DOF = 1 satisfies both requirements.
Proof Sketch. Case analysis on $\mathbb{N}$: For DOF = 1, any two queries return the single location’s value; transitivity of equality forces agreement. For DOF $\geq$ 2, construct two locations with values $v$ and $v^{\prime}\neq v$; queries return different answers. This witness construction works uniformly for all DOF $\geq$ 2. By trichotomy of naturals, DOF = 1 is the unique solution.
We introduce the zero-incoherence capacity: the maximum rate guaranteeing zero disagreement among replicated encodings. Main results: exact capacity ($C_{0}=1$), tight side-information bound ($\geq\log_{2}k$ bits for $k$-way incoherence), and rate-complexity separation ($O(1)$ at capacity vs $\Omega(n)$ above).
Encoding locations are terminals in multi-terminal source coding. Derivation is perfect correlation reducing effective rate; only complete derivation achieves zero incoherence. We give achievability and converse proofs, formalize confusability/incoherence graphs, and present the mutual-information side-information bound.
Constructive instantiations (programming patterns, software case study) supplement theory. Extended evaluation, code, and Lean proofs (6300+ lines, 0 sorry) are in supplementary material.
Index Terms: zero-error capacity, multi-terminal source coding, side information, confusability graph, impossibility theorems
@@ I Introduction
I-A Single-Source Coherence Theorem
We begin with a proof from first principles. Consider any system that encodes a fact at $n$ independent locations. We prove that exactly one value of $n$ guarantees coherence: $n=1$.
Theorem (Single-Source Coherence). Among all possible degrees of freedom (DOF = number of independent encoding locations), exactly one value (DOF = 1) guarantees coherence. DOF = 0 fails (no fact encoded). DOF $\geq$ 2 fails (permits explicit construction of inconsistency).
Proof Sketch. Case analysis on $\mathbb{N}$: For DOF = 1, any two queries return the single location’s value; transitivity of equality forces agreement. For DOF $\geq$ 2, construct two locations with values $v$ and $v^{\prime}\neq v$; queries return different answers. By trichotomy of naturals, DOF = 1 is the unique solution (TRI1). (Formalized in Lean 4, 0 sorry.)
This theorem is the foundation. It establishes that single-source-of-truth (SSOT) is not merely preferable but mathematically forced for systems requiring coherence (UNQ1). The rest of this paper derives the information-theoretic consequences.
I-B Zero-Incoherence Capacity
Building on the Single-Source Coherence Theorem, we ask: what encoding rate guarantees exactly zero probability of disagreement among replicated encodings? An encoding system stores a fact $F$ at locations $\{L_{1},\dots,L_{n}\}$. The system is coherent when all locations agree. The central quantity is the zero-incoherence capacity $C_{0}$, the supremal rate achieving exactly zero incoherence probability. Our capacity theorem follows directly:
@@ I-C The Capacity Theorem
I-D Applications Across Domains
@@
I-E Connection to Classical Information Theory
@@ I-F Encoder Realizability
I-G Paper Organization
Core theorems (capacity, realizability, complexity bounds) are machine-checked in Lean 4 [15] (3758 lines, 185 theorem/lemma statements, 0 sorry placeholders). The entropy and mutual-information arguments in Section VI-D are encoded as explicit classical-information assumptions in the Lean closure layer, matching paper-level conditional usage of Fano-style bounds [13].
@@ I-H Core Theorems
Uniqueness. $C_{0}=1$ is the unique capacity: DOF =0=0 fails to encode; DOF >1>1 exceeds capacity. Given zero-incoherence as a constraint, the rate is mathematically forced (UNQ1, MDL5).
I-I Scope
@@ 1 Coherence guaranteed (optimal rate under coherence constraint)
This uniqueness is forced by trichotomy (TRI1): the three cases are exhaustive, and only DOF = 1 achieves coherence (UNQ1). This is formalized as the Zero-Incoherence Capacity theorem (Theorem II.36) in Section II-K.
(L: INS1)
@@ III-D Derivation: The Coherence Mechanism
Structural facts require definition-time derivation. Structural facts (class existence, schema structure, service topology) are fixed when defined. Compile-time derivation that runs before the definition is fixed is too early (the declarative source is not yet fixed). Runtime is too late (structure already immutable). Definition-time is the unique opportunity for structural derivation—a consequence of timing trichotomy (TRI1, MDL3, REQ1, REQ2).
@@ IV-H The Logical Chain (Summary)
Mechanization status. The core requirement chain is machine-checked in Lean 4; current build statistics are 3758 lines, 185 theorem/lemma statements, and 0 sorry placeholders.